# Navigating the Latent Space Dynamics of Neural Models

### v2

@@ Abstract
Neural Network (NN) transform high-dimensional data into compact, structured representations, often modeled as elements of a lower dimensional latent space. In this paper, we present an alternative interpretation of neural models as dynamical systems acting on the latent manifold. Specifically, we show that AutoEncoder (AE) models implicitly define a latent vector field on the manifold, derived by iteratively applying the encoding-decoding map, without any additional training. We observe that standard training procedures introduce inductive biases that lead to the emergence of attractor points within this vector field. Drawing on this insight, we propose to leverage the vector field as a representation for the network, providing a novel tool to analyze the properties of the model and the data. This representation enables to: $(i)$ analyze the generalization and memorization regimes of neural models, even throughout training; $(ii)$ extract prior knowledge encoded in the network’s parameters from the attractors, without requiring any input data; $(iii)$ identify out-of-distribution samples from their trajectories in the vector field. We further validate our approach on vision foundation models, showcasing the applicability and effectiveness of our method in real-world scenarios.
@@ 1 Introduction
Figure 1: The latent dynamics of AEs Latent vector fields with $\mathbf{z}_{0}\sim\mathcal{U}[-8,8]$ of 2D AEs trained on the MNIST dataset with different initializations. Colors denote the norm of the vectors: the boundary denotes the shape of the latent manifold identified with the support of the encoder, in white we observe regions of the latent space where the vector field vanishes, flowing toward attractors, which correspond to regions of high probability density of the data.
Neural models are effective function approximators that have been demonstrated to be able to solve complex tasks in countless domains [23, 41]. These models typically learn to transform high-dimensional data into compact, structured representations [4], often modeled as elements of a lower dimensional latent space. In this work, we propose a novel alternative representation for neural models by showing that, for a given AE architecture, there exists an associated latent vector field (see Figure 1), which characterizes the behavior of the network at the representation level. The vector field is obtained by simply iterating the encoding decoding map in the latent space, without any additional training procedures. Previous work [42, 22] has demonstrated that in the extreme overparametrized regime and under strong assumptions, AEs can effectively memorize training data. In this work, we propose to consider the more general case in which memorization corresponds to one extremum, showing evidence that the latent vector field induced by general AEs can be reconnected to properties of the network, such as generalization, memorization, and out-of-distribution detection. Our contributions can be summarized as follows:
  • We show that every AE architecture induces a latent vector field, implicitly. We demonstrate that this representation retains properties of the networks and the data, throughout its trajectories and its fixed points.
  • We show that most mappings learned by NNs are contractive: this enforces the presence of fixed points and attractors points in the latent vector field.
  • We empirically show how attractors characterize the memorization and generalization regime in NNs and their interplay in the training process.
  • We experimentally demonstrate the information stored in vision foundation models weights can be retrieved in a data-free way, by probing models with noise and retrieve attractors.
  • Finally, we show that paths in the latent vector fields are informative to characterize the learned distribution and to detect distribution shifts.
@@ 2 Method
In the following sections we will consider neural models $F$ as a composition of encoder decoder modules $F_{\Theta}=D_{\theta_{2}}\circ E_{\theta_{1}}$ parametrized by $\Theta=[\theta_{1},\theta_{2}]$. The encoder maps samples $\mathbf{x}\sim p(\mathbf{x})$ supported on $\mathcal{X}\subset\mathbb{R}^{m}$ to a typically lower dimensional space $\mathcal{Z}\subset\mathbb{R}^{k}$ and the network $F$ is trained for reconstruction to minimize the mean squared error $\mathcal{L}_{MSE}$:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in X}\|\mathbf{x}-F_{\Theta}(\mathbf{x})\|_{2}^{2}+\lambda\mathcal{R}(\Theta)$ (1)
where $\mathcal{R}$ is either an explicit (e.g., weight decay) or implicit regularization constraint on the parameters $\Theta$ to favor contractive solutions of $F$. We posit that minimizers of the above objective act as minimizers of the spectral norm of the Jacobian $\rho(J_{F}(x))$ for $x\sim p(x)$. Some examples of regularizer include weight decay, where $\mathcal{R}=\|\Theta\|_{2}^{F}$ or the use of some augmentations distribution $p(T)$, for which the objectives becomes:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in X}\|\mathbf{x}-F(\mathbf{x})\|_{2}^{2}+\sum_{T\in p(T)}\|\mathbf{x}-F(T\mathbf{x})\|_{2}^{2}$ (2)
examples of $p(T)$ include additive gaussian noise in denoising AEs [52] or masking, in masked AEs [15]. The bottleneck dimension $k=dim(\mathcal{Z})$ is another example of how to enforce a contractive behavior in the network. More precisely, the bottleneck dimension acts as a hard constraint on the rank of the Jacobian of the encoder $rank(J_{E})<=k$ and therefore of the entire map $F$, since $J_{F}(\cdot)=J_{E}(\cdot)\odot J_{D}(\cdot)$. In Table C in the Appendix, we list many AE variants including denoising AEs (DAEs) [52], sparse AEs (SAEs) [37], variational AEs (VAEs)  [26] and other variants [44, 2, 12] and show how their objectives enforce local contractive solutions around training points.
Given an AE model $F$ possibly pretrained, denoting with $f(\mathbf{z})=E\circ D(\mathbf{z})$, we are interested in analyzing the iterations of the map $f(\ldots f(f(\mathbf{z})))$, which can be modeled as a differential equation.
@@ 2.1 The Latent Dynamics of Neural Models
For a sample $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{k}$ we analyze the effect of applying iteratively $f$ on $\mathbf{\mathbf{z}}$, i.e. $f\circ f\circ\ldots f(\mathbf{z})$. Parametrizing the number of iteration by $t$, this lead to the following discrete ordinary differential equation (ODE):
$\begin{cases}&\mathbf{z}_{t+1}=f(\mathbf{z})\\ &\mathbf{z}_{0}=\mathbf{z}\end{cases}$ (3)
@@
In Eq 3, one can associate the map $f$ as the pushforward of a latent vector field $V:\mathbb{R}^{k}\mapsto\mathbb{R}^{k}$ tracing non linear trajectories in the latent space (see Fig 1). A first interesting question is whether this ODE has solutions and if they are unique. Through the Banach fixed point theorem, it can be shown that the ODE has a unique solution iff the map $f$ is Lipschitz continuous.
@@
A function $f:\mathcal{Z}\mapsto\mathcal{Z}$ is denoted Lipschitz continuous if there exist a constant $C$ s.t. for every pair of points $\mathbf{x_{1}},\mathbf{x_{2}}$:
$d(\mathbf{z_{1}},\mathbf{z_{2}})_{\mathcal{X}}\leq C\ d(f(\mathbf{z_{1}}),f(\mathbf{z_{2}}))_{\mathcal{X}}$ (5)
when $C<1$ $f$ is called contractive.
For any contractive map, there will exist fixed point solutions of Eq 3, i.e., iteratively applying the map $f$ will converge to a unique constant solution via solving $\mathbf{z}^{*}=f(\mathbf{z}^{*})$. The fixed point $\mathbf{x}$ solutions correspond to singular points that attract and summarize the dynamics of the system.
@@
A fixed point $\mathbf{x}=f(\mathbf{x})$ is an attractor of a differentiable map $f$ if all eigenvalues of the Jacobian $J$ of $f$ at $\mathbf{x}$ are strictly less than one in absolute value.
When $f$ is nonlinear, the ODE in Eq. 3 can have multiple solutions depending on the initial conditions $\mathbf{z_{0}}$. The previous definition allows for a definition of Lipschitz continuity which is inherently local:
@@
[…]
@@
The set of initial conditions $Z_{0}$ leading to the same attractor $\mathbf{z}^{*}$, is denoted as basin of attraction.
Why maps learned by NNs are contractive in practice? We argue that most maps learned by neural AEs or variants thereof, promote (locally) contractive autoencoding mappings. In other words, many explicit or implicit regularization strategies and training design choices constrain the Jacobian of the autoencoding map to have small eigenvalues near training examples. We provide below what are the inductive biases that promote locally contractive mappings:
  • Initialization Standard initialization schemes [29, 14, 16] aim at preserving the variance of the activations at initialization to avoid gradient vanishing and exploding problems during training. Theoretical arguments in support assume i.i.d. input samples and weights, and are generally tied to specific activation functions (e.g., ReLU in [16]). We note that in practice, this condition is not likely to hold, as training data samples are often correlated and many employed architectural elements (for example, residual connections [17]) break the independence assumptions of the layer weights. Networks result in a bias to be either expansive or contractive towards a single point, with a preference for the latter case [39]. We show empirically that this is indeed the case in Figure 6 in the Appendix, across several vision backbone models. In Figure 3, we show an example of a 2D latent vector field which is globally contractive at initialization (Figure 3(a), left plot) towards a single attractor at the origin. This is the case even in higher dimensions (see the number of attractors at the first epoch in Figure 3(c)).
  • Explicit regularization Widely employed regularization techniques such as weight decay [7], implicitly constraints the learned map to be (locally) contractive. This can be seen easily in the case of linear mappings, where penalizing the norm of the matrix corresponds to penalizing the first-order derivatives of the map, resulting in the optimal way to make the map contractive. Although this is not guaranteed to be optimal for nonlinear networks, weight decay still enforces a contraction mapping. The regularizer is incorporated in state-of-the-art optimizers [32], ubiquitous in state-of-the-art foundation models. Other forms of parametric constraints include the choice of bottleneck dimension, which acts as a hard constraint on the rank of $J_{F}(\mathbf{z})$, and soft variants thereof, such as the Kullback-Leibler divergence term in VAEs, or sparsity regularization in SAEs.
  • Implicit regularization As previously discussed, the choice of augmentation distribution implicitly selects the notion of neighborhood of the training samples (for example, Gaussian radius in DAEs), and the perturbation direction that will be penalized in $J_{f}(\mathbf{x})$. This form of regularization is nonparametric and inherently local as opposed to the explicit regularization techniques explained in the above paragraph.
@@ 3 Theoretical remarks
In this section, we are interested in analyzing properties of trajectories within the latent vector field, as well as its fixed points and attractors, when they exist.
@@ 3.1 Trajectories of the latent vector field
In [2, 34, 45], it was shown that, under optimal conditions, the residual in denoising AEs trained with noise variance $\sigma^{2}$ approximates the score function, i.e., the gradient of the log empirical density of the data, as $\sigma\rightarrow 0$. Interestingly, [2] shows a connection between denoising and contractive AEs, where the norm of the Jacobian of the input-output mapping is penalized. The more general result we want to stress here is that whenever the auto-encoding map is locally contractive w.r.t. a given notion of neighborood (e.g. gaussian radius, masking augmentations) and approximates well enough the input distribution $p(\mathbf{x})$, then the associated latent vector field pushes samples in the direction of the derivative of the score of the approximated prior in the latent space. Informally, this means that the vector field will non-linearly project samples on the data manifold onto regions of high probability density.
Theorem 1.
[…]
In general, attractors cannot be computed via solving the fixed point equation with first-order methods, unless the initial condition is arbitrary close to the attractors, as the following result shows:
Proposition 3.1.
Iterations of the map $f$ in Equation 3 correspond to performing gradient descent in the $\frac{\partial L(\mathbf{z})}{{\partial\mathbf{z}}}$ iff $f$ is an isometry (i.e. its eigenvalues are concentrated near 1) or in the proximity of attractor points, when the Jacobian of $f$ vanishes.
This result highlights the importance of the higher order terms involved in iterating $f$, which trace a highly nonlinear trajectory in latent space, escaping local minima solutions given by unstable and spurious fixed points.

### v3

@@ Abstract
Neural networks transform high-dimensional data into compact, structured representations, often modeled as elements of a lower dimensional latent space. In this paper, we present an alternative interpretation of neural models as dynamical systems acting on the latent manifold. Specifically, we show that autoencoder models implicitly define a latent vector field on the manifold, derived by iteratively applying the encoding-decoding map, without any additional training. We observe that standard training procedures introduce inductive biases that lead to the emergence of attractor points within this vector field. Drawing on this insight, we propose to leverage the vector field as a representation for the network, providing a novel tool to analyze the properties of the model and the data. This representation enables to: $(i)$ analyze the generalization and memorization regimes of neural models, even throughout training; $(ii)$ extract prior knowledge encoded in the network’s parameters from the attractors, without requiring any input data; $(iii)$ identify out-of-distribution samples from their trajectories in the vector field. We further validate our approach on vision foundation models, showcasing the applicability and effectiveness of our method in real-world scenarios.
@@ 1 Introduction
Figure 1: Latent dynamics of AEs. Latent vector fields induced by autoencoders with bottleneck $k=2$, trained on MNIST, with $\mathbf{z}_{0}\sim\mathcal{U}[-8,8]$. Models with different initializations are shown. Colors (viridis colormap) represent vector norms ranging from violet (low) to yellow (high). The shape of the latent manifold identifies with the encoder’s support. White regions indicate where the vector field vanishes, revealing attractors aligned with high-density areas of the data distribution.
Neural networks are powerful function approximators, capable of solving complex tasks across a wide range of domains (Jumper et al., 2021; Radford et al., 2021). A core component of their success lies in their ability to transform high-dimensional inputs into compact, structured representations (Bengio et al., 2013), typically modeled as points in a lower-dimensional latent space.
In this work, we propose a novel perspective: for any given autoencoder (AE) architecture, there exists an associated latent vector field (see Figure 1) that characterizes the model’s behavior in representation space. This field arises naturally by iterating the encoder-decoder map in the latent space, requiring no additional training. Intuitively, such dynamics often settle into attractors: stable states toward which nearby trajectories converge, summarizing the long-term behavior of the system.
While prior work has shown that overparameterized autoencoders can memorize training data under strong assumptions (Radhakrishnan et al., 2020; Jiang and Pehlevan, 2020), we consider a broader setting. We show that memorization emerges as just one type of attractor in the latent vector field, and that the structure of this field reflects deeper properties of the network, such as generalization behavior and sensitivity to distribution shifts.
Our contributions can be summarized as follows:
  • We show that every AE implicitly defines a latent vector field, whose trajectories and fixed points encode properties of both the model and the data.
  • We demonstrate that most neural mappings are contractive, leading naturally to the emergence of fixed points and attractors in the latent space.
  • We empirically connect attractors to the network’s memorization and generalization regimes, showing how they evolve during training.
  • We show experimentally that vision foundation models can be probed in a data-free manner: initializing with noise, we recover attractors that reveal semantic information embedded in the model’s weights.
  • Finally, we show that trajectories in the latent vector field can capture the learned data distribution and serve as a signal for detecting distribution shifts.
@@ 2 Method
We consider neural models $F$ as compositions of encoder-decoder modules, defined as $F_{\Theta}=D_{\theta_{2}}\circ E_{\theta_{1}}$ parametrized by $\Theta=[\theta_{1},\theta_{2}]$. The encoder $E_{\theta_{1}}$ maps inputs $\mathbf{x}\sim p(\mathbf{x})$ supported on $\mathcal{X}\subset\mathbb{R}^{m}$ to a typically lower-dimensional space $\mathcal{Z}\subset\mathbb{R}^{k}$, and the decoder $D_{\theta_{2}}$ reconstructs the input. The model is trained to minimize the mean squared error $\mathcal{L}_{MSE}$:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in X}\|\mathbf{x}-F_{\Theta}(\mathbf{x})\|_{2}^{2}+\lambda\mathcal{R}(\Theta)\,$, (1)
where $\mathcal{R}$ is either an explicit or implicit regularization term encouraging contractive behavior in $F$.
We posit that minimizing this objective leads to a reduction in the spectral norm of the Jacobian, $\|J_{F}(\mathbf{x})\|_{\sigma}$ for $\mathbf{x}\sim p(\mathbf{x})$. Examples of regularization include weight decay, where $\mathcal{R}=\|\Theta\|_{2}^{F}$, or data augmentation by sampling transformations $T\sim p(T)$ and minimizing:
$\mathcal{L}_{MSE}(\mathbf{x})=\sum_{\mathbf{x}\in\mathcal{X}}\|\mathbf{x}-F(\mathbf{x})\|_{2}^{2}+\sum_{T\in p(T)}\|\mathbf{x}-F(T\mathbf{x})\|_{2}^{2}\,.$ (2)
Such augmentations include additive Gaussian noise in denoising AEs (Vincent et al., 2008) and input masking in masked AEs (He et al., 2022). Another form of contractive pressure is the bottleneck dimension $k=\dim(\mathcal{Z})$, which places a hard upper bound on the rank of the encoder Jacobian, $\mathrm{rank}(J_{E})<=k$ and, by the chain rule, constrains the full model Jacobian $J_{F}(\cdot)=J_{E}(\cdot)\odot J_{D}(\cdot)$. Table C (Appendix) lists many AE variants, including denoising AEs (DAEs) (Vincent et al., 2008), sparse AEs (SAEs) (Ng et al., 2011), variational AEs (VAEs) (Kingma et al., 2013), and others (Rifai et al., 2011; Alain and Bengio, 2014; Gao et al., 2024), highlighting how their objectives promote local contractive behavior around training data.
Given a possibly pretrained AE model, we define the map $f(\mathbf{z})=E\circ D(\mathbf{z})$ and study its repeated application $f(\ldots f(f(\mathbf{z})))$, which can be modeled as a differential equation.
@@ 2.1 The Latent Dynamics of Neural Models
Given a sample $\mathbf{z}\in\mathcal{Z}\subseteq\mathbb{R}^{k}$, we study the effect of repeatedly applying the map $f$, i.e., $f\circ f\circ\ldots f(\mathbf{z})$. By introducing a discrete time parameter $t$, this iterative process defines the discrete ODE:
$\begin{cases}&\mathbf{z}_{t+1}=f(\mathbf{z}_{t})\\ &\mathbf{z}_{0}=\mathbf{z}\end{cases}$ (3)
@@
In Eq. 3, the map $f$ defines the pushforward of a latent vector field $V:\mathbb{R}^{k}\mapsto\mathbb{R}^{k}$, tracing nonlinear trajectories in latent space (Fig. 1). A natural question is whether the ODE has well-defined and unique solutions. By the Banach fixed-point theorem, this holds if and only if $f$ is Lipschitz-continuous.
@@
A function $f:\mathcal{Z}\mapsto\mathcal{Z}$ is Lipschitz-continuous if there exists a constant $C$ s.t. for every pair of points $\mathbf{z_{1}},\mathbf{z_{2}}$:
$d(f(\mathbf{z_{1}}),f(\mathbf{z_{2}}))_{\mathcal{Z}}\leq C\ d(\mathbf{z_{1}},\mathbf{z_{2}})_{\mathcal{Z}}\,.$ (5)
When $C<1$, $f$ is called contractive, for a given metric $d$ on $\mathcal{Z}$.
For any contractive map, Eq. 3 admits fixed-point solutions, i.e., repeatedly applying the map $f$ will converge to a unique solution $\mathbf{z}^{*}$ satisfying $\mathbf{z}^{*}=f(\mathbf{z}^{*})$. The fixed points $\mathbf{z}^{*}$ can act as attractors, capturing and summarizing the system’s long-term dynamics.
@@
A fixed point $\mathbf{z}=f(\mathbf{z})$ is an attractor of a differentiable map $f$ if all eigenvalues of the Jacobian $J$ of $f$ at $\mathbf{z}$ are strictly less than one in absolute value.
When $f$ is nonlinear, the ODE in Eq. 3 can have multiple solutions depending on the initial conditions $\mathbf{z_{0}}$. The previous definition allows for a definition of Lipschitz continuity, which is inherently local:
@@
Let $f$ : $\mathcal{Z}\mapsto\mathcal{Z}$ be differentiable and $C$-Lipschitz continuous. Then the Lipschitz constant $C$ is given by: $C$ = $sup_{\mathbf{z}\in\mathcal{Z}}\|J_{f}(\mathbf{z})f\|_{\sigma}$ , where $J_{f}(\mathbf{z})$ is the Jacobian of $f$ evaluated at $\mathbf{z}$ and $\|\cdot\|_{\sigma}$ is the spectral norm.
@@
The set of initial conditions $\mathbf{z}_{0}$ leading to the same attractor $\mathbf{z}^{*}$, is denoted as basin of attraction.
Why are neural mappings contractive in practice? We argue that mappings learned by neural AEs and their variants tend to be locally contractive, i.e.,their Jacobians have small eigenvalues near training examples. This behavior emerges naturally from several explicit and implicit inductive biases present in modern training pipelines. Below, we outline the main factors promoting contractive behavior:
  • Initialization bias. Standard initialization schemes (LeCun et al., 2002; Glorot and Bengio, 2010; He et al., 2015) are designed to preserve activation variance at the start of training to avoid vanishing or exploding gradients. Theoretical arguments in support assume i.i.d. inputs and weights, and are often tied to specific activation functions (e.g., ReLU in (He et al., 2015)). However, real-world training data is typically correlated, and architectural features like residual connections (He et al., 2016) break weight independence. As a result, networks often exhibit a bias toward mappings that are globally expansive or contractive, empirically skewed toward the latter (Poole et al., 2016).
    We illustrate this empirically in Figure 6 (Appendix), showing that various vision backbones exhibit contractive behavior at initialization. Figure 3(a) (left) shows a 2D latent vector field which is globally contractive at initialization towards a single attractor at the origin. This contractive behavior also holds in higher dimensions (see the number of attractors at epoch 0 in Figure 3(c)).
  • Explicit regularization. Common regularization methods like weight decay (D’Angelo et al., 2024) encourage contraction by penalizing the norm of model parameters. In the linear case, minimizing parameter norms directly reduces the Jacobian’s spectral norm, making the map contractive. While this link is less direct for nonlinear models, the effect due to weight decay persists in practice. Regularization is often integrated into optimizers used in large-scale models (Loshchilov, 2017). Additional architectural constraints, such as small bottleneck dimensions, also limit the rank of $J_{F}(\mathbf{z})$. Soft constraints include KL divergence in VAEs or sparsity penalties in SAEs (see Table C in the Appendix for a complete overview).
  • Implicit regularization. Data augmentations introduce local perturbations around training examples, effectively defining a neighborhood structure in input space. For instance, Gaussian noise in DAEs or masking in MAEs perturbs inputs along specific directions. These augmentations implicitly regularize the Jacobian $J_{f}(\mathbf{x})$ by penalizing sensitivity to those perturbations. Unlike parametric regularization, this effect is inherently local and nonparametric.
@@ 3 Theoretical remarks
In this section, we study the behavior of the latent vector field, focusing on its trajectories (Section 3.1) as well as its fixed points and attractors (Section 3.2), when they exist. Formal statements and proofs of all theorems and propositions are provided in Appendix A.
@@ 3.1 Trajectories of the latent vector field
Prior work (Miyasawa et al., 1961; Robbins, 1992; Alain and Bengio, 2014) has shown that, under ideal conditions, the residual of a denoising AE trained with noise variance $\sigma^{2}$ approximates the score function, i.e., the gradient of the log-density of the data, as $\sigma\rightarrow 0$. Interestingly, Alain and Bengio (2014) shows a connection between denoising and contractive AEs, where the Jacobian norm of the input-output mapping is explicitly penalized.
We build on this by observing a more general phenomenon: when an autoencoding map is locally contractive relative to a chosen neighborhood structure (e.g., Gaussian noise, masking) and sufficiently approximates the input distribution $p(\mathbf{x})$, the induced latent vector field pushes points in the direction of the score of the corresponding prior in latent space.
Informally, this implies that the vector field acts to nonlinearly project samples toward regions of high probability on the data manifold.
Theorem 1 (informal, proof in Appendix A.1).
Let $F$ be a trained autoencoder and let $q(\mathbf{z})=\int p(\mathbf{x})q(\mathbf{z}|\mathbf{x})d\mathbf{x}$ be the marginal distribution induced in latent space. Assume $q(\mathbf{z})$ is smooth and that there exists an open neighborhood $\Omega\supseteq\mathrm{supp}\,q$ and a constant $L<1$ such that $\sup_{\mathbf{z}\in\Omega}\bigl\lVert J_{f}(z)\bigr\rVert_{\sigma}\;\leq\;L$. Then, latent dynamics $f(\mathbf{z})-\mathbf{z}$ in $\Omega$ is proportional to the score function $\nabla_{\mathbf{z}}\log q(\mathbf{z})$.
This result establishes a general link between the vector field’s trajectories and the score function, under the assumption of local contractivity. In practice, neural networks often strike a balance between the reconstruction loss (Eq. 1) and regularization on the Jacobian, which contributes to the emergence of attractors in the latent space.
An important implication of Theorem 1 is that integrating the vector field $f(\mathbf{z})-\mathbf{z}$ effectively estimates the log-density, i.e., $\int_{\mathbf{z}}\nabla\log q(\mathbf{z})dz=\log q(\mathbf{z})+C$ in unbounded domains. If the Jacobian $J_{f}$ is symmetric (e.g., in AEs with tied weights (Alain and Olivier, 2013)), then the latent vector field is conservative and corresponds to the gradient of a potential $V_{f}=\nabla E$. In this case, the AE defines an energy-based model with $q(\mathbf{z})=e^{-E(\mathbf{z})}$ (Zhai et al., 2016; Song and Kingma, 2021). In the general (non-conservative) case, we show empirically in Section 4.2 that the trajectories still reflect the learned prior distribution.
However, attractors are generally not accessible solving the fixed-point equation with first order methods, unless initialized very close to them. The following result formalizes this:
Proposition 3.1 (informal, proof in Appendix A.2).
Iterations of the map $f$ in Eq. 3 correspond to gradient descent on a potential $L(\mathbf{z})$ (i.e., $f\approx\mathbf{z}-\alpha L(\mathbf{z})$) if and only if $f$ is an isometry (its eigenvalues are near 1), or the dynamics occur near attractors, where $J_{f}$ vanishes.
This underscores the role of higher-order dynamics in the vector field: repeated applications of $f$ trace complex, nonlinear trajectories that can escape spurious or unstable fixed points.